# Can Transformers Do Enumerative Geometry?

### v1

@@
Abstract.
How can Transformers model and learn enumerative geometry? What is a robust procedure for using Transformers in abductive knowledge discovery within a mathematician-machine collaboration? In this work, we introduce a new paradigm in computational enumerative geometry in analyzing the $\psi$-class intersection numbers on the moduli space of curves. By formulating the enumerative problem as a continuous optimization task, we develop a Transformer-based model for computing $\psi$-class intersection numbers based on the underlying quantum Airy structure. For a finite range of genera, our model is capable of regressing intersection numbers that span an extremely wide range of values, from $10^{-45}$ to $10^{45}$. To provide a proper inductive bias for capturing the recursive behavior of intersection numbers, we propose a new activation function, Dynamic Range Activator (DRA). Moreover, given the severe heteroscedasticity of $\psi$-class intersections and the required precision, we quantify the uncertainty of the predictions using Conformal Prediction with a dynamic sliding window that is aware of the number of marked points. Next, we go beyond merely computing intersection numbers and explore the enumerative “world-model” of the Transformers. Through a series of causal inference and correlational interpretability analyses, we demonstrate that Transformers are actually modeling Virasoro constraints in a purely data-driven manner. Additionally, we provide evidence for the comprehension of several values appearing in the large genus asymptotic of $\psi$-class intersection numbers through abductive hypothesis testing.
Figure 1. The Dilaton relations matrix (left) for $\psi$-class intersections at $g=14$. Cosine similarity $S_{i,j}$ between learned embeddings (hidden representation) of predicted intersection numbers (right) and a filtered similarity matrix with a threshold of $S_{i,j}\geq 0.98$ (middle). This provides evidence that the Transformer has (re)discovered and internalized the Dilaton equation.
1. Introduction
Enumerative geometry is a fascinating branch of mathematics concerned with counting geometric objects or finding invariants that satisfy certain geometric conditions (see for instance [mumford1983towards, katz2006enumerative]). The objects of interest are typically complex curves that need to match a specific geometric blueprint. A classical question in enumerative geometry is: How many conics pass through five given points in the plane? A more contemporary question is: How many rational curves of degree one are there on a quintic threefold?
Approaching such problems often requires recursive computational methods. However, these methods can become computationally expensive and even infeasible, even for relatively small values of the parameters of interest. This is where we propose using Machine Learning (ML), particularly Transformers [vaswani2023attentionneed], to overcome this challenge. While recent work with Transformers, especially in Large Language Models (LLMs) [zhao2023survey], has focused on their ability to solve problems in natural language processing and image generation, these models still struggle with complex recursive and mathematical reasoning tasks in a zero-shot setting. This limitation is particularly critical in fields that require deep, structured mathematical reasoning. By incorporating Transformers as a primary tool in enumerative geometry, specifically for approximating $\psi$-class intersection numbers, we aim to address the computational complexity overhead and explore several important questions. Traditional methods for computing these numbers, given genus $g$ and $n$ marked points, are computationally expensive with a complexity of $O(g!\cdot 2^{n})$. These intersection numbers are famously computed by the Korteweg–de Vries equation [witten1991two, kontsevich1992intersection] and are of paramount importance in the study of various geometric and physical problems, such as Weil–Petersson volumes, Hurwitz numbers, Gromov–Witten invariants, and topological 2D quantum gravity. Going further, we investigate whether our proposed model is capable of causal and abductive reasoning for knowledge discovery. Specifically, we ask if we can conduct parametric inference and extract mathematical insights from its internal understanding, beyond the raw data provided, to offer intuition and potentially generate new conjectures.
@@ 1. Introduction
It is well known that Transformer-based models have difficulty with even periodic patterns, let alone recursive problems [ziyin2020neural]. Therefore, we first examine the architectural design choices and training pipeline variants necessary for learning the recursive reasoning processes involved in topological recursion [eynard2007invariants]. To this end, we introduce a non-linear activation function, the Dynamic Range Activator (DRA), which is particularly suited for learning recursive functions. Consequently, we develop DynamicFormer, a modified Transformer-based model designed to predict $\psi$-class intersection numbers given the quantum Airy structure [kontsevich2018airy] initial data. Additionally, we believe that a predictive model without uncertainty estimation is inherently unreliable. Therefore, we incorporate uncertainty quantification into our model’s predictions using Conformal Prediction [shafer2008tutorial].
To the best of our knowledge, this is the first time Transformers have been applied to enumerative problems in algebraic geometry, while also employing interpretability analysis [chefer2021transformer, liu2024interpretable, kantamneni2024transformers] to extract knowledge and illuminate the model’s internal understanding of deep research-level mathematical concepts. Thus, in this paper, we aim to target both machine learning scientists and mathematicians, as our work is relevant to both communities.
Figure 2. Illustration of DynamicFormer. It processes two main input modalities: the quantum Airy structure $B$ as an ordered sequence and partitions $d$ as a permutation-invariant set. The genus $g$ and number of marked points $n$ are incorporated as input properties, modulated with the main modalities at various stages. All layers, including the Multi-Head Attention (MHA) blocks, use the Dynamic Range Activator (DRA) non-linear activation function. The MHA block for $B$-modality integrates Dynamic Positional bias with linear positioning as well. The Talking-Modalities block consists of Batch Normalization [ioffe2015batch] and the self-supervised loss. The DRA prediction head is a 2-layer MLP that predicts $\psi$-class intersection numbers in logarithmic scale.
2. Related Works
Transformer-based models [vaswani2023attentionneed], which were originally developed for natural language processing [radford2018improving, devlin2019bert], have since been leveraged for various data modalities [dosovitskiy2021, gong2021astaudio, carolan2024reviewmultimodal]. Due to their flexibility and interoperability, they have been extensively used in different tasks in Artificial Intelligence (AI) for Science [wang2023scientific]. In recent years, a new branch of ML applications, often referred to as AI for Mathematics [he2023machine], has emerged. This branch includes Automated Theorem Proving [loos2017deepnetworkguidedproof, paliwal2020graph, an2024learnfailurefinetuningllms], Generation [trinh2024solving, wang2020learning, lin2024atg], Autoformalization [wang2020exploration, wu2022autoformalization, jiang2023multilingual, mikula2024magnushammer, ying2024lean], machine-human collaborative guidance [bao2020quiver, anderson2021moduli, davies2021advancing, berman2022machine, craven2023learning, coates2023machine, dong2024machine], examples/counterexamples search problems [halverson2019branes, gukov2021learning, wagner2021constructions, gukov2023searching, berglund2024generating, romera2024mathematical], and predictive problems (such as amplitude and mathematical expression prediction) [saxton2019, dersy2023simplifying, alnuqaydan2023symba, dascoli2022deep, cai2024transforming, coates2024machine]. These areas extensively use Transformers, particularly in problems arising from particle interactions and scattering processes.
@@ 2. Related Works
Our work bridges the domains of machine-human collaborative guidance and predictive problems while introducing new methods. In the prediction task, while most studies remain within in-distribution settings, we extend our approach to out-of-distribution predictions, coupled with Conformal Prediction uncertainty estimation—a notably absent requirement in current literature. Additionally, we enhance the predictive task by using multi-modal data, involving both discrete sets and continuous graph (sequence) representations. This approach moves beyond the focus on mere compact symbolic/numerical sequence modalities [meidani2024snip]. Furthermore, in the field of machine-human collaborative guidance, we conduct causal inference and correlational analyses to understand the “world-model” of Transformers [micheli2023transformers], offering deep insights into the underlying algebraic geometry problem at hand.
3. Mathematical Background
In this section, we review some background on the original mathematical problem. Let ¯$\overline{\mathcal{M}}_{g,n}$ be the moduli space of genus $g\geq 0$ complex stable algebraic curves with $n>0$ distinct marked points. This is a complex, compact, smooth orbifold of dimension $3g-3+n$. Associated with each marked point is a natural cohomology class $\psi_{i}=c_{1}(\mathcal{L}_{i})$, defined as the first Chern class of the cotangent line bundle $\mathcal{L}_{i}$ at the $i$-th marked point.
In the early 90s, Witten [witten1991two] analyzed a theory of two-dimensional topological quantum gravity, where $\psi$-classes play the role of observables, and the intersection numbers (sometimes called amplitudes)
$\langle\tau_{d_{1}}\cdots\tau_{d_{n}}\rangle_{g,n}\coloneqq\int_{\overline{\mathcal{M}}_{g,n}}\psi_{1}^{d_{1}}\cdots\psi_{n}^{d_{n}}\,,\qquad d_{1}+\cdots+d_{n}=d_{g,n}\coloneqq 3g-3+n\,,$ (3.1)
represent correlators of the theory, and are topological invariants of ¯$\overline{\mathcal{M}}_{g,n}$. Based on a low genus profile and considerations from physics, Witten conjectured that the exponential generating function
$Z(\bm{x};\hbar)=\exp\left(\sum_{\begin{subarray}{c}g\geq 0\\ n>0\end{subarray}}\frac{\hbar^{2g-2+n}}{n!}\sum_{\begin{subarray}{c}d_{1},\ldots,d_{n}\geq 0\\ d_{1}+\cdots+d_{n}=d_{g,n}\end{subarray}}\left\langle\tau_{d_{1}}\cdots\tau_{d_{n}}\right\rangle_{g,n}x^{d_{1}}\cdots x^{d_{n}}\right)$ (3.2)
satisfies a infinite tower of quadratic partial differential equations with respect to the formal variables $(x^{i})_{i\geq 0}$ known as the Korteweg–de Vries (KdV) hierarchy. This was later proven by Kontsevich [kontsevich1992intersection].
In recent years, it has been discovered that many fundamental invariants in physics and geometry can be described via intersection numbers. In physics, interest in $\psi$-class intersection numbers has been revitalized due to their connection with Jackiw–Teitelboim gravity [saad2019jt]. From an algebraic geometry perspective, $\psi$-class intersection numbers represent the most basic form of Gromov–Witten invariants [behrend1997gromov]. These intersection numbers are central to enumerative problems addressed by [kontsevich1992intersection] and [mirzakhani2007weil] regarding the symplectic volumes of the moduli space of metric ribbon graphs and hyperbolic Riemann surfaces. Computing intersection numbers is pivotal in the asymptotic counting of geodesic curves in both flat and hyperbolic random geometries [mirzakhani2008growth, delecroix2021masur, andersen2023topological].
Given the wide and fascinating applications of intersection numbers, computing them for arbitrarily large genera remains an open problem. Their association with the KdV hierarchy has been illuminating for calculating these invariants [witten1991two, kontsevich1992intersection]. Another approach to characterizing these numbers involves Virasoro constraints [Dijkgraaf:1990dj], where the partition function is annihilated by a collection of differential operators that form a representation of the Virasoro algebra. Alternative recursive approaches include topological recursion [eynard2007invariants], the cut-and-join equation [alexandrov2011cut], and exact formulae for the generating series, such as error-function-type integrals [okounkov2002generating] and the determinantal formula [bergere2009determinantal]. However, the general computation of intersection numbers remains an open problem, as these methods can only be used to compute $\psi$-class intersection numbers for cases with small $g$ or $n$. Meanwhile, universality phenomena in flat and hyperbolic geometry and in the dynamics of surfaces manifest themselves in large genera. Only recently, Aggarwal [aggarwal_large] has found a closed-form formula for the large genus asymptotic of $\psi$-class intersection numbers.
3.1. Quantum Airy Structures
@@
The conceptually simplest approach to computing $\psi$-class intersection numbers involves using Virasoro constraints recast in the language of quantum Airy structures. The latter is an algebraic reformulation of topological recursion of Eynard and Orantin [eynard2007invariants] due to Kontsevich and Soibelman [kontsevich2018airy]. A quantum Airy structure on a (possibly countably infinite-dimensional) complex vector space $V$, dually generated by $(x^{i})_{i\in I}$, is a family of differential operators $(L_{i})_{i\in I}$ in the Weyl algebra over $V$ of the form
@@
and forming a Lie sub-algebra of the Weyl algebra:
$[L_{i},L_{j}]=\hbar^{2}\sum_{a\in I}f^{a}_{i,j}L_{a}\,,\qquad\qquad\forall i,j\in I\,.$ (3.4)
Here $\partial_{i}=\partial_{x^{i}}$ and $A_{i,a,b}=A_{i,b,a}$, $B_{i,a}^{b}$, $C_{i}^{a,b}=C_{i}^{b,a}$, $D_{i}$, and $f^{a}_{i,j}=-f^{a}_{j,i}$ are scalars indexed by $i,j,k\in I$. We remark that $I$ is the set of indices used to label the basis elements of $V$. In the infinite-dimensional setting, certain boundedness conditions must be imposed to guarantee that all sums are finite [andersen2024abcd].
Given a quantum Airy structure, one can uniquely find a suitably normalized formal function $Z$ on $V$, called partition function, which is annihilated by the differential operators $(L_{i})_{i\in I}$. More precisely, the following theorem holds.
@@
The elements $F_{g,n}=\sum_{d_{1},\ldots,d_{n}\in I}F_{g;d_{1},\ldots,d_{n}}\,x^{d_{1}}\cdots x^{d_{n}}$ are symmetric tensors of rank $n$ over $V^{*}$. The collection of coefficients $(A,B,C,D)$ can also be regarded as tensors, and we will often refer to them as the quantum Airy structure data.
For a given quantum Airy structure, the coefficients $F_{g;d_{1},\ldots,d_{n}}$, called quantum Airy structure amplitudes, are determined recursively on (minus) the Euler characteristic $2g-2+n$. For the lowest values of $g$ and $n$, the base cases are given by
$F_{0;i,j,k}\coloneqq A_{i,j,k}\,,\qquad\qquad F_{1;i}\coloneqq D_{i}\,,$ (3.7)
where $A_{i,j,k}$ and $D_{i}$ are part of the data of the quantum Airy structure. For higher values of $g$ and $n$ satisfying $02g-2+n>$ 0, the recursion formula, known as topological recursion, is given by
@@
[…]
The simplest example of a quantum Airy structure involves taking $V=\mathbb{C}$ and setting $A=B=C=D=1$. Under these conditions, we have a unique differential operator given by
$L=\hbar\partial_{x}-\hbar^{2}\left(\frac{x^{2}}{2}+x\partial_{x}+\frac{1}{2}\partial_{x}^{2}+1\right).$ (3.9)
The differential equation $LZ=0$ can be reduced to the Airy differential equation (hence the name “quantum Airy structure”). Therefore, the associated formal power series $Z$ is given by the properly normalized asymptotic expansion of the Airy function:
$Z(x;\hbar)=\frac{e^{\frac{x}{\hbar}(1-\frac{\hbar x}{2})}}{\mathrm{Bi}(0)}\mathrm{Bi}\Biggl{(}\frac{1-2\hbar x-\hbar^{2}}{(2\hbar^{2})^{2/3}}\Biggr{)}.$ (3.10)
@@
For higher-dimensional vector spaces $V$, finding a closed-form solution becomes increasingly impractical, if not impossible. Besides, the most interesting applications of quantum Airy structures involve an infinite-dimensional vector space $V$. The central example is the one associated to the $\psi$-class intersection numbers. In this case, the underlying vector space is generated by vectors indexed by $I=\mathbb{N}$, the non-negative integers. The differential operators $(L_{i})_{i\geq 0}$ are determined by the tensors
$A_{i,j,k}\coloneqq\delta_{i,j,k}\,,$ (3.11)
$B^{k}_{i,j}\coloneqq\delta_{i+j,k+1}\frac{(2k+1)!!}{(2i+1)!!(2j-1)!!}\,,$
$C^{j,k}_{i}\coloneqq\delta_{i,j+k+2}\frac{(2j+1)!!(2k+1)!!}{(2i+1)!!}\,,$
$D_{i}\coloneqq\frac{\delta_{i,1}}{24}\,.$
@@
The resulting operators, after shifting of the indices as $\mathsf{L}_{i}=L_{i-1}$, form a representation of the Virasoro algebra: $[\mathsf{L}_{i},\mathsf{L}_{j}]=\hbar^{2}(i-j)\mathsf{L}_{i+j}$ for all $i\geq-1$. Remarkably, the associated amplitudes coincide with $\psi$-class intersection numbers:
@@
$F_{g;d_{1},\ldots,d_{n}}=\begin{cases}\langle\tau_{d_{1}}\cdots\tau_{d_{n}}\rangle_{g,n}&\text{if }d_{1}+\cdots+d_{n}=d_{g,n}\,,\\ 0&\text{otherwise.}\end{cases}$ (3.12)
@@
In this paper, we focus on the specific choice of quantum Airy structure data given by equation 3.11, thereby computing the $\psi$-class intersection numbers. For ease of notation, we will denote a generic partition of $d_{g,n}$ of length $n$ as $\bm{d}=(d_{1},\ldots,d_{n})$. The associated $\psi$-class intersection number will be denoted by
$\langle\bm{d}\rangle_{g,n}\coloneqq\langle\tau_{d_{1}}\cdots\tau_{d_{n}}\rangle_{g,n}.$ (3.13)
The goal of the next sections is to provide evidence for the possibility of regressing these numbers via a Transformer-based model, namely DynamicFormer, given the quantum Airy structure initial data. Our model is trained on known data computed by a brute force algorithm up to genus $13$ and is tested up to genus $16$. We then provide causal and correlational evidence that our model is actually learning and understanding the recursive nature of the intersection numbers. As a result, we aim to introduce a new paradigm for using Transformers in algebraic geometry as both a powerful and explainable amortization method and a source of intuition for the mathematician.
4. Data Representation and Setup
    $\bm{B}=\Set{(i,j,k,B_{i,j}^{k})}{B_{i,j}^{k}\neq 0}$ (4.1)
    $\bm{C}=\Set{(i,j,k,C_{i}^{j,k})}{C_{i}^{j,k}\neq 0}$ (4.2)
    where $i,j,k\in[0,d_{g,n}]$, and $C_{i}^{j,k}$ are the non-zero components of the tensor. With this choice of $d_{\textup{max}}$, the maximum length of $\bm{C}$ is approximately 1500.
@@ 4. Data Representation and Setup
We trained the model using data up to genus $g=13$ and then tasked it with predicting the intersection numbers for genera $g=14$, $15$, and $16$. During experimentation, we observed that, when comparing the contributions of $B$ and $C$, the $B$ tensor had a greater impact on $\psi$-class intersection numbers. Consequently, we incorporated only this initial data (excluding $C$), a decision further supported by the inherent properties of the intersection numbers. This initial observation motivated the subsequent series of explainability analysis experiments and findings.
5. Methods
To handle the different input modalities, i.e. the continuous tensor $B$ in COO format and the discrete set $\bm{d}\in\mathbb{N}^{n}$, we developed a Transformer-based model as depicted in figure 2. However, managing different input formats is only part of the challenge. The recursive nature of many functions in enumerative geometry adds further complexity. A key property of recursive maps, such as the factorial function, is their rapid growth. Learning this recursive behavior requires not only fitting patterns within a bounded region but also successfully extrapolating those patterns beyond that region. However, it has been shown that Transformers are not very successful at capturing recursive behavior [lakretz2021causal, zhang2024transformerbased]. To address the limitations of Transformers in learning the recursive behavior of invariants in enumerative geometry, we introduce DynamicFormer to capture such behavior and perform proper inference for recursive problems.
Dynamic Positional Bias. The trunk of the network consists of two modified Transformers: one discrete, acting on the embedding of the partitions $\bm{d}$ and maintaining permutation equivariance, and one continuous, acting on the $B$ tensor [vaswani2023attentionneed]. The positions of elements in $B$, represented in COO format, correspond to particular combinations of indices that compute the intersection numbers. To enhance generalization for longer input sequences, we integrated a modified version of Dynamic Positional Bias [wang2021crossformer, zhang2024transformerbased]. Dynamic Positional Bias computes a relative positional bias map for each attention head, which is learnable and adapts based on sequence length. However, because it does not inherently account for the distance from the start or end of a COO sequence, we apply masking during training based on input sequence length to address this limitation.

### v3

@@
Abstract
We introduce a Transformer-based approach to computational enumerative geometry, specifically targeting the computation of $\psi$-class intersection numbers on the moduli space of curves. Traditional methods for calculating these numbers suffer from factorial computational complexity, making them impractical to use. By reformulating the problem as a continuous optimization task, we compute intersection numbers across a wide value range from $10^{-45}$ to $10^{45}$. To capture the recursive nature inherent in these intersection numbers, we propose the Dynamic Range Activator (DRA)  [^1] GitHub Code: https://github.com/Baran-phys/DynamicFormer, a new activation function that enhances the Transformer’s ability to model recursive patterns and handle severe heteroscedasticity. Given the precision required to compute these invariants, we quantify the uncertainty in the predictions using Conformal Prediction with a dynamic sliding window, adaptive to partitions of equivalent numbers of marked points. To the best of our knowledge, there has been no prior work on modeling recursive functions with such a high-variance and factorial growth. Beyond simply computing intersection numbers, we explore the enumerative “world-model” of Transformers. Our interpretability analysis reveals that the network is implicitly modeling the Virasoro constraints in a purely data-driven manner. Moreover, through abductive hypothesis testing, probing, and causal inference, we uncover evidence of an emergent internal representation of the the large-genus asymptotic of $\psi$-class intersection numbers. These findings suggest that the network internalizes the parameters of the asymptotic closed-form and the polynomiality phenomenon of $\psi$-class intersection numbers in a non-linear manner.
1 Introduction
Enumerative geometry is a branch of mathematics concerned with counting geometric objects or finding invariants that satisfy certain geometric conditions (Mumford, 1983; Katz, 2006). A classical question in enumerative geometry is: How many conics pass through five given points in the plane? A more contemporary one is: How many rational curves of degree one are there on a quintic threefold? Addressing such problems often boils down to computing intersection numbers, which are fundamental in computing Weil–Petersson volumes, Hurwitz numbers, Gromov–Witten invariants, and quantum gravity amplitudes. An example of such objects is $\psi$-class intersection numbers. Traditional recursive methods like the KdV hierarchy (Witten, 1991; Kontsevich, 1992) have computational complexities that grow like $O(g!\cdot 2^{n})$, where $g$ is the genus and $n$ is the number of marked points, making calculations infeasible even for low values of $g$ and $n$. To overcome this computational bottleneck, we propose employing Machine Learning (ML), particularly Transformers (Vaswani et al., 2017), to approximate $\psi$-class intersection numbers.
@@ 1 Introduction
It is well-known that Transformer-based models have difficulty with even periodic patterns, let alone recursive problems (Ziyin et al., 2020). Therefore, we first examine the architectural design choices and training pipeline variants necessary for learning the recursive reasoning processes involved in quantum Airy structures (Kontsevich & Soibelman, 2018; Andersen et al., 2024). Hence, we introduce a non-linear activation function, the Dynamic Range Activator (DRA), which is particularly suited for learning recursive functions. Consequently, we develop DynamicFormer, a modified Transformer-based model designed to predict $\psi$-class intersection numbers given the quantum Airy structure initial data. Additionally, we believe that a predictive model without uncertainty estimation is inherently unreliable. Therefore, we incorporate uncertainty quantification into our model’s predictions using Conformal Prediction (Shafer & Vovk, 2008). Going further, we investigate whether the network is capable of abductive reasoning for knowledge discovery. Specifically, we ask whether we can conduct parametric inference and extract mathematical insights from its internal understanding beyond the raw data provided to offer intuition and potentially generate new conjectures. We provide causal and correlational evidence that the network is actually understanding the underlying mathematics. As a result, we aim to introduce a new paradigm for using Transformers in algebraic geometry as both a powerful and explainable amortization method and a source of intuition for the mathematician.
While the studies discussed in Appendix B primarily focus on symbolic logic or compact numerical sequences, they fall short in handling the highly recursive and complex structures found in enumerative geometry. Existing models are limited to in-distribution settings and struggle with the high-variance and recursive nature of problems in computational algebraic geometry. For example, Belcak et al. (2022) discuss and benchmark various ML models only up to unique-type functions and do not even enter the realm of recursive functions.
@@
Our work bridges machine-human collaborative guidance and predictive problems while introducing new methods. Most studies in prediction tasks remain within in-distribution and compact settings, whereas we extend our approach to out-of-distribution recursive predictions, incorporating Conformal Prediction uncertainty estimation—an often overlooked aspect in the literature. Precision is essential in mathematics, especially when a mathematician needs to assess whether an informed conjecture is correct based on available samples. In such decision-making processes, having access to the uncertainty of AI-predicted samples/values is crucial, as it provides insight into the reliability of the predictions and guides the validation of mathematical hypotheses. We also enhance the predictive task by using multi-modal data, involving both discrete sets and continuous graph (sequence) representations. This approach moves beyond the focus on mere compact symbolic/numerical sequence modalities (Meidani et al., 2024). Furthermore, in the field of machine-human collaborative guidance, we conduct causal inference and correlational analyses to understand the mathematical “world-model” of Transformers (Li et al., 2023; Micheli et al., 2023), offering deep insights into the underlying algebraic geometry problem at hand.
To the best of our knowledge, this is the first time Transformers and explainable ML have been applied to enumerative algebraic geometry to extract knowledge and illuminate the model’s internal understanding of deep research-level mathematical concepts. In the end, we aim to target both machine learning scientists and mathematicians, as our work is relevant to both communities.
2 Quantum Airy Structures
@@
The conceptually simplest approach to computing $\psi$-class intersection numbers involves using Virasoro constraints recast in the language of quantum Airy structures (Kontsevich & Soibelman, 2018; Andersen et al., 2024). The latter is an algebraic reformulation of topological recursion (Eynard & Orantin, 2007). A quantum Airy structure on a complex vector space $V$, dually generated by $(x^{i})_{i\in I}$, is a family of differential operators $(L_{i})_{i\in I}$ in the Weyl algebra over $V$ of the form
@@
and forming a Lie sub-algebra of the Weyl algebra. Here $\partial_{i}=\partial_{x^{i}}$ and $A_{i,a,b}=A_{i,b,a}$, $B_{i,a}^{b}$, $C_{i}^{a,b}=C_{i}^{b,a}$, and $D_{i}$, are scalars indexed by elements in $I$, and $\hbar$ is a formal parameter that keeps track of the genus. Given a quantum Airy structure, one can uniquely find a formal function $Z$ on $V$, called partition function, which is annihilated by the differential operators $(L_{i})_{i\in I}$. More precisely, the following theorem holds.
@@
The elements $F_{g,n}=\sum_{d_{1},\ldots,d_{n}\in I}F_{g;d_{1},\ldots,d_{n}}\,x^{d_{1}}\cdots x^{d_{n}}$ are symmetric tensors of rank $n$ over $V^{*}$. The collection of coefficients $(A,B,C,D)$ can also be regarded as tensors, and we will often refer to them as the quantum Airy structure initial data. For a given quantum Airy structure, the coefficients $F_{g;d_{1},\ldots,d_{n}}$, called quantum Airy structure amplitudes, are determined recursively on (minus) the Euler characteristic $2g-2+n$. For the lowest values of $g$ and $n$, the base cases are given by $F_{0;i,j,k}\coloneqq A_{i,j,k}$ and $F_{1;i}\coloneqq D_{i}$, where $A_{i,j,k}$ and $D_{i}$ are part of the data of the quantum Airy structure. For higher values of $g$ and $n$ satisfying $02g-2+n>$ 0, the recursion formula, known as topological recursion, is given by
@@
In general, the computation of the amplitudes has a computational complexity of $O(g!\cdot 2^{n})$, which makes the calculation of the amplitudes at high genera problematic. For high-dimensional vector spaces $V$, finding a closed-form solution becomes increasingly impractical, if not impossible. The central example is the one associated to the $\psi$-class intersection numbers. In this case, the underlying vector space is generated by vectors indexed by $I=\mathbb{N}$, the non-negative integers. The differential operators $(L_{i})_{i\geq 0}$ are determined by the tensors
$A_{i,j,k}\coloneqq\delta_{i,j,k}\,,\quad$ $B^{k}_{i,j}\coloneqq\delta_{i+j,k+1}\frac{(2k+1)!!}{(2i+1)!!(2j-1)!!}\,,$ (2.5)
$C^{j,k}_{i}\coloneqq\delta_{i,j+k+2}\frac{(2j+1)!!(2k+1)!!}{(2i+1)!!}\,,\quad$ $D_{i}\coloneqq\frac{\delta_{i,1}}{24}\,.$
@@
The resulting operators, after shifting of the indices and rescaling as $\mathsf{L}_{i}=-\frac{(2i-1)!!}{2}L_{i-1}$, form a representation of the Virasoro algebra: $[\mathsf{L}_{i},\mathsf{L}_{j}]=\hbar^{2}(i-j)\mathsf{L}_{i+j}$ for all $i\geq-1$. Remarkably, the associated amplitudes coincide with $\psi$-class intersection numbers (see Appendix C):
@@
$F_{g;d_{1},\ldots,d_{n}}=\begin{cases}\langle\tau_{d_{1}}\cdots\tau_{d_{n}}\rangle_{g,n}&\text{if }d_{1}+\cdots+d_{n}=d_{g,n}\coloneqq 3g-3+n\,,\\ 0&\text{otherwise.}\end{cases}$ (2.6)
@@
In this paper, we focus on the specific choice of quantum Airy structure data given by Equation 2.5, thereby computing the $\psi$-class intersection numbers. For ease of notation, we will denote a generic partition of $d_{g,n}$ of length $n$ as $\bm{d}=(d_{1},\ldots,d_{n})$. The associated $\psi$-class intersection number will be denoted by $\langle\bm{d}\rangle_{g,n}\coloneqq\langle\tau_{d_{1}}\cdots\tau_{d_{n}}\rangle_{g,n}$.
@@
The goal of the next sections is to provide the possibility of regressing these numbers via a Transformer-based model, given the quantum Airy structure initial data. Our model is trained on known data computed by a brute force algorithm up to genus $13$ and is tested up to genus $17$. During experimentation, we observed that when comparing the contributions of $B$ and $C$, the $B$ tensor had a greater impact on $\psi$-class intersection numbers. Consequently, we incorporated only this initial datum (excluding $C$), a decision further supported by the inherent properties of the intersection numbers (see  Appendix D). This initial observation motivated the subsequent series of explainability analysis experiments and findings.
3 Methods
The recursive nature of functions in enumerative geometry, combined with the sparse and high-variance target distributions of $\psi$-class intersection numbers, introduces a high complexity in modeling and accurately approximating these functions. A main property of recursive maps (e.g factorial function), is their dramatic growth and drop. Learning this recursive behavior requires not only fitting high-frequency patterns within a bounded region but also successfully extrapolating those patterns beyond that region. In time series prediction tasks, capturing periodic even behavior is a challenge. Various methods (Miller et al., 2024) have been employed to model periodic patterns effectively. However, these approaches typically deal with uni-modal data that also exhibit relatively low variance in both In-Distribution (ID) and Out-Of-Distribution (OOD) regions and do not generalize well to recursive problems with the high-variance observed in our context. Therefore, traditional methods for modeling high-variance recursive data, including those used in time series analysis and standard Transformer architectures, are insufficient for our needs (Lakretz et al., 2021; Belcak et al., 2022; Zhang et al., 2024). Thus, to capture such behavior and perform proper inference for multi-modal recursive problems, we enhance Transformers by introducing the Dynamic Range Activator (DRA), and introduce DynamicFormer, depicted in Figure 5. The DRA is designed to handle the recursive and factorial growth properties inherent in enumerative problems with minimal computational overhead and can be integrated into existing neural networks without requiring significant architectural changes.